# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ 4. Any critical point of $\mathcal{L}$ is a KKT point of problem (24);
3. Since the sequence $\{(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})\}_{t\geq 1}$ is bounded, there exists a convergent subsequence $\{(\mathbf{z}^{t_{k}},\mathbf{w}^{t_{k}},\bm{\lambda}^{t_{k}})\}_{k\geq 1}$ that converges to a limit point $(\mathbf{z}^{*},\mathbf{w}^{*},\bm{\lambda}^{*})$. As $t\to\infty$, combining Lemma A.8.1 and equation (B), we obtain
@@
On the other hand, by Lemma A.8.1, equation (B), and Lemma A.8.3, we know that $\mathbf{w}^{t_{k}}\in\partial\mathcal{P}^{t_{k}}$ and $\mathbf{w}^{t_{k}}\to 0$ as $k\to\infty$. Since the subdifferential $\partial\mathcal{P}$ is closed see ([[35], Remark 1(ii)), we have
@@
Furthermore, from equation (G), it follows that
@@
Combining equations (G) and (G) shows that $(\mathbf{z}^{*},\mathbf{w}^{*},\bm{\lambda}^{*})$ is a critical point of $\mathcal{L}$.
@@
$\mathbf{0}$ $\in\partial_{\mathbf{w}_{i}}\mathcal{L}_{i}(\mathbf{z}_{i}^{*},\mathbf{w}_{i}^{*},\bm{\lambda}_{i}^{*})=\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{*},\mathbf{w}_{i}^{*})+\delta_{\mathcal{B}^{N_{i}+1}}(\mathbf{w}_{i}^{*}),$ (G.25)
@@
Since for any fixed $\mathbf{w}_{i}$, the function $G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})+\langle\bm{\lambda}_{i},\mathbf{A}_{i}\mathbf{z}_{i}\rangle$ is convex in $\mathbf{z}_{i}$ over the set $\mathcal{X}\cap\mathcal{Y}$, equation (G) implies that $\mathbf{z}_{i}^{*}$ also satisfies
@@
Combining (G.25), (G.26), and (G), we conclude that $(\mathbf{z}^{*},\mathbf{w}^{*},\bm{\lambda}^{*})$ is a KKT point of problem (24).
@@
1. There exists an integer $\overline{t}\geq 0$ such that $\mathcal{P}^{\overline{t}}=\mathcal{P}^{*}$. Then from Lemma A.8.1 and equation (G), we have
@@
2. For all $t$, we have $\mathcal{P}^{t}>\mathcal{P}^{*}$. From equation (G), for any $\eta>0$, there exists $t_{0}>0$ such that
@@
Combining this with equations (G) and (G), and the fact that $\varphi^{\prime}(\mathcal{P}^{t}-\mathcal{P}^{*})>0$, yields:
@@
Substituting equation (G) into (G) gives
@@
Summing equation (G) with $k=t_{2}+1,t_{2}+2,\cdots,t$, we obtain

### v2

@@ 4. Any critical point of $\mathcal{L}$ is a KKT point of problem (23);
3. Since the sequence $\{(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})\}_{t\geq 1}$ is bounded, there exists a convergent subsequence $\{(\mathbf{z}^{t_{k}},\mathbf{w}^{t_{k}},\bm{\lambda}^{t_{k}})\}_{k\geq 1}$ that converges to a limit point $(\mathbf{z}^{*},\mathbf{w}^{*},\bm{\lambda}^{*})$. As $t\to\infty$, combining Lemma A.8.1 and equation (A.2), we obtain
@@
On the other hand, by Lemma A.8.1, equation (A.2), and Lemma A.8.3, we know that $\mathbf{w}^{t_{k}}\in\partial\mathcal{P}^{t_{k}}$ and $\mathbf{w}^{t_{k}}\to 0$ as $k\to\infty$. Since the subdifferential $\partial\mathcal{P}$ is closed see ([[32], Remark 1(ii)), we have
@@
Furthermore, from equation (A.7), it follows that
@@
Combining equations (A.7) and (A.7) shows that $(\mathbf{z}^{*},\mathbf{w}^{*},\bm{\lambda}^{*})$ is a critical point of $\mathcal{L}$.
@@
$\mathbf{0}$ $\in\partial_{\mathbf{w}_{i}}\mathcal{L}_{i}(\mathbf{z}_{i}^{*},\mathbf{w}_{i}^{*},\bm{\lambda}_{i}^{*})=\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{*},\mathbf{w}_{i}^{*})+\delta_{\mathcal{B}^{{\color[rgb]{0,0,0}N_{i}+M}}}(\mathbf{w}_{i}^{*}),$ (G.25)
@@
Since for any fixed $\mathbf{w}_{i}$, the function $G_{i}(\mathbf{z}_{i},\mathbf{w}_{i})+\langle\bm{\lambda}_{i},\mathbf{A}_{i}\mathbf{z}_{i}\rangle$ is convex in $\mathbf{z}_{i}$ over the set $\mathcal{X}\cap\mathcal{Y}$, equation (A.7) implies that $\mathbf{z}_{i}^{*}$ also satisfies
@@
Combining (G.25), (G.26), and (A.7), we conclude that $(\mathbf{z}^{*},\mathbf{w}^{*},\bm{\lambda}^{*})$ is a KKT point of problem (23).
@@
1. There exists an integer $\overline{t}\geq 0$ such that $\mathcal{P}^{\overline{t}}=\mathcal{P}^{*}$. Then from Lemma A.8.1 and equation (A.7), we have
@@
2. For all $t$, we have $\mathcal{P}^{t}>\mathcal{P}^{*}$. From equation (A.7), for any $\eta>0$, there exists $t_{0}>0$ such that
@@
Combining this with equations (A.7) and (A.7), and the fact that $\varphi^{\prime}(\mathcal{P}^{t}-\mathcal{P}^{*})>0$, yields:
@@
Substituting equation (A.7) into (A.7) gives
@@
Summing equation (A.7) with $k=t_{2}+1,t_{2}+2,\cdots,t$, we obtain